### v2

# Explaining Context Length Scaling and Bounds for Language Models

Abstract
Long Context Language Models have drawn great attention in the past few years. There has been work discussing the impact of long context on Language Model performance: some find that long irrelevant context could harm performance, while some experimentally summarize loss reduction by relevant long context as Scaling Laws. This calls for a more thorough understanding on how long context impact Language Modeling. In this work, we (1) propose a clean and effective theoretical framework on explaining the impact of context length to Language Modeling, from an Intrinsic Space perspective; and (2) conduct experiments on natural language and synthetic data, validating our proposed theoretical assumptions and deductions. Our theoretical framework can provide practical insights such as establishing that training dataset size dictates an optimal context length and bounds context length scaling for certain case. We hope our work may inspire new long context Language Models, as well as future work studying Physics for Language Models. Code for our experiments is available at this anonymous url: https://anonymous.4open.science/r/xxxxxx.
Deep Learning, Context Length, Language Modeling
1 Introduction
Because of the rapid development of capacity of Language Models and the importance of a long context length in tasks like reasoning, retrieval, etc, past years people have been attempting to extend the context length of Language Models. There have been a variety of methods on supporting long context Language Models(Su et al., 2023; Katharopoulos et al., 2020; Gu & Dao, 2024; Peng et al., 2023; Sun et al., 2024). A wide variety of work is proposed to discuss the impact of context length: some shows long irrelevant context would worsen performance for LMs(Xu et al., 2024; Levy et al., 2024); some shows long context would improve performance in a way summarized as Scaling Laws(Xiong et al., 2024); while work in other domains like time series shows long relevant context would hurt performance (Shi et al., 2024). This calls for a more thorough understanding of how context length affects Language Models’ performance..
Previously, theories have been proposed to explain the Scaling Laws with respect to the data set and the size of the model(Bahri et al., 2024; Sharma & Kaplan, 2020). However, these theories do not study how context length impact Language Modeling, thus they cannot contribute directly to the problem.
In this work, we propose a theory framework to discuss the impact of context length from an Intrinsic Space perspective. Starting with simple assumptions w.r.t. Intrinsic Space and Intrinsic Dimension, we come up with a clean and effective theoretical explanation of relationship between Cross Entropy Loss, Intrinsic Dimension and Context Length. We further use real language and synthetic data to validate our assumptions and deductions. Our main contributions include,
  • 1. We propose a theoretical framework on understanding Language Modeling for different context length in Language Models from the perspective of Intrinsic Space.
  • 2. We conduct experiments on real and synthetic data, validating our theoretical assumptions and deductions.
Our theoretical framework can naturally derive phenomena that we have observed. For example, our theory derives that, for a certain amount of training data, as the context length increases, the neural network first would first behave like the Bayes Model (thus the loss decreases); then beyond a certain optimal context length, the gap between the trained model and the Bayes Model would increase, hence the validation loss would increase: this is experimentally verified, as shown in Figure 1,
We hope our work may inspire future work when it comes to explaining context impact and/or designing new long context Language Models. For example, we observe that longer contexts have less impact (or there is a ceiling on the information about the next token as context length approaches infinity), discouraging models with excessively large context lengths. On the other hand, these findings have the potential to provide guarantees for RNNs with finite hidden states in long-sequence modeling.
Figure 1: Upper figure: Validation Loss vs. Context Length, measured on subsets of OpenWebText dataset. We see for each training dataset size, there exists an optimal context length that minimizes pretraining validation loss, which increases with the dataset size (More details can be found in Section 3). Lower figure: similar results obtained on our Synthetic Dataset in Section 4.5. This proves the deduction of our theory.
Our phenomenological theory in the main paper (mainly presented in Section 2) based on Intrinsic Dimension is simpler and more intuitive, while it can also be explained by a more fundamental theory based on Entropy, as presented in Appendix C.
2 Assumptions, Deductions and Observations for Language Modeling
2.1 Preliminary: Loss Decomposition and Intrinsic Space
It is common in ML studies to decompose the loss into Bayes Risk (the minimum loss possible, achieved by the theoretically optimal Bayesian Model), and Approximation Loss (the loss measuring the ability of a trained model actually to approximate the Bayesian Model). Specifically for Cross-Entropy loss, we have (please refer to Appendix A for more details):
$H(P,Q_{l})=$ $R_{Bayes}+L_{Approximate}$ (1)
$=$ $H(P,P_{l})+L_{Approximate}(P_{l},Q_{l})$
Where $P=p(x_{0}|x_{-\infty:0})$ is the distribution of Natural Language (or our experimented dataset), $P_{l}=p(x_{0}|x_{-l:0})$ is the Bayesian Model for context length $l$ and $Q=q(x_{0}|x_{-l:0})$ is the learned Language Model. $H(P,P_{l})$ is the Bayes Risk of optimal model (the assumed ‘limit’ when we have infinite data points and model parameters) and $L_{Approximate}(D,N,l,\ldots)$ is the Approximation Loss.
Bayesian Models are typically related to the Intrinsic Space of data manifold in such a way: the middle-layer or last-layer data representation of a well-trained Bayesian Neural Network is often used to approximate the Intrinsic Space(Cheng et al., 2023). In Section 2.2 we discuss more how this bridges Bayes Risk with Intrinsic Space, and how context length impacts Bayes Risk.
Approximation Loss, or how well the trained model learns Bayesian Model, is also related to Intrinsic Dimension in the perspective of Scaling Laws. In Section 2.3 we discuss more about how the context length impacts Approximation Loss from this perspective.
We use a synthetic data set to prove concepts in Section 4.
We further derive that the balance between Bayes Risk and Approximation Loss would lead to an optimal context length which increases with the size of the training dataset. Our theoretical deduction and experiments on language and synthetic data are presented in Section 3 and Section 4.5.
2.2 Bayes Risk with context length: an intrinsic space perspective
In this section we discuss the impact of context length on the Bayes Risk, from an Intrinsic Space and Intrinsic Dimension perspective; we also conduct experiments measuring the dependency of context length to Bayes Risk.
2.2.1 Bayes Risk and Entropy in Intrinsic Space: derived from First Principles
We propose a simple theory model to relate $H(P,P_{l})$ with the intrinsic dimension $dim(l)$ of the intrinsic space $\text{space}_{l}$ of the text corpora of length $l$ (for the next-token prediction task).
From first-principles we assume that,
  • Assumption 1. Intrinsic Dimension of the Bayes Model $\lim_{l\rightarrow\infty}dim(l)=dim(\infty)$ is finite, which is the Intrinsic Dimension of next token prediction of language itself.
  • Assumption 2. $\forall l_{1},l_{2}\text{ such that }l_{1}<l_{2}$, $dim(l_{1})<dim(l_{2})$. This is because a longer context contains more information about the next possible token.
To simplify deduction, we further assume that,
  • Assumption 3. Each intrinsic dimension would add $s$ bits of information to the next-token prediction task, so there are $dim(l)*s$ bits of information that can be represented in $\text{space}_{l}$ for the next-token prediction. Note this does not mean these are the only information in the Intrinsic Space, hence $s$ can be small, or even smaller than $1$.
  • Assumption 4. KL-divergence for the Bayes ¡+Model of context length $l$, $P_{l}$, with Bayes Model of infinite context length, $P=P_{\infty}$, equal to $s*(dim(\infty)-dim(l))$.
We further assume that in intrinsic space the states are with equal probability. With these assumptions, we can derive $H(P,P_{l})$ with $dim(l)$ (please refer to Appendix A for more details on the definition of Cross-Entropy loss used in this work, and Appendix B for a detailed derivation):
$R_{Bayes}(l)$ $=H(P,P_{l})$ (2)
$=H(P)+D_{KL}(P,P_{l})$
$\approx H(P)+s*(dim(\infty)-dim(l))$
$=-s*dim(l)+H(P)+s*dim(\infty).$
This linear relationship is observed in experiments for LMs in Section 2.2.2, and for synthetic data in Section 4.3.
Note that by Assumptions 1 and 2 we derive that:
$\frac{\partial R_{Bayes}}{\partial l}<0\text{, and }\lim_{l\rightarrow\infty}\frac{\partial R_{Bayes}}{\partial l}=0.$ (3)
2.2.2 Bayes Risk and Intrinsic Dimension: Experiment Measurement
We use well-trained Large Language Models to conduct experiments for approximating the Bayes Risk $H(P_{l})$ on certain text corpora.
Figure 2: Bayes Risk vs. Context Length: Bayes Risk is approximated by Cross Entropy loss measured with LLaMa-3.1 series on OpenWebText, for different context length.
We find that:
$H(P_{l})\approx C_{0}+C/l^{\gamma}$ (4)
approximates the experimented behavior on OpenWebText well. Please see Figure 2 for the result. Moreover, we further conduct experiments on a dataset that is sured not to be included in LLaMa 3.1 8B’s pretraining dataset. Please see further information in Appendix D.
We further use PCA as a metric to measure the Intrinsic Dimension of Dataset with respect to context length. Figure 3 provides the relative degradation of the eigenvalue in the feature space of LLaMa-3.1-8B, for the last token. We see that larger input length would indeed provide feature with lower degradation in the intrinsic space. Notably, when $15005<idx<$ 1500 the curves is similar to Zip-f distribution ($\log eig=C_{0}-C*\log idx$), and for $4000500<idx<$ 4000 it resembles exponential degradation ($\log eig=C_{0}-C*idx$). Naturally, w ‘Intrinsic Dimension’. However, these formulas are very similar around $idx\approx 1000$, hence cannot provide accurate estimations to the transformation index.
Instead, following previous practice, here we use some threshold to decide the Ine would like to use the transformation index of these two states as Itrinsic Dimension: $\max_{idx}rela\_eig(idx)\geq\text{threshold}$ is used as the measured Intrinsic Dimensionz. We further evaluate the validity of this method in Point 2 in Section 4.1, on a synthetic Dataset with synthetic data; and explain the linear measurement from an Entropy perspective in Appendix C. Notably, the threshold here is a hyperparameter which is set to constants like $1/20$ in some previous work(Aghajanyan et al., 2020), but we observe here that many thresholds would validate the linear correspondence of Cross Entropy vs. Intrinsic Dimension, which further enhance the robustness of our result. In detail, as shown in Figure 3 and Figure 4, we use thresholds from $0.002$ to $0.25$ to make our conclusion more robust.
For a certain threshold, we conduct experiments on several context lengths, and measure CE Loss on certain text corpora with these context lengths. We observe a fairly linear relationship between CE Loss and ID measured (supporting our theory), as shown in Figure 4:
Figure 3: Relative Eigen Value for LLaMa-3.1-8B on a subset of OpenWebText, presented in different x-axis scales, with different context length visible to Language Model. Gray lines represent different thresholds we take to measure the intrinsic dimension of the current model.
Figure 4: Cross Entropy Loss vs. Measured Intrinsic Dimension, for LLaMa-3.1-8B on a subset of OpenWebText. Each line represents a certain threshold used to measure ID in the intrinsic space of the used LLM. Different Measurements would give ID values that are linear w.r.t. each other, and they are all linear w.r.t. CE loss.
We see from Figure 4 that, no matter what threshold we use, the Cross Entropy Loss usually follows a linear relationship with the Intrinsic Dimension we measured, showing the robustness of the PCA evaluation method, and validating our theoretical assumptions:
$dim(l)\approx-s*dim(l)+Const$
Which aligns well with Equation 2, thus validating our previous deduction.
We use the concept of Intrinsic Dimension and Measured Intrinsic Dimension to simplify the deduction. Moreover, by using the perspective of $S=\log\Omega$ where $\Omega$ is the number of states in Intrinsic Space, similar results can be derived, and why the measured IDs are linear w.r.t. each other and CE loss can be (partially) explained. Please refer to Appendix C for further discussions.
2.3 Approximation loss with Context Length: an intrinsic dimension perspective
In previous work people have experimentally summarize the Scaling Laws (Kaplan et al., 2020; Hoffmann et al., 2022) as: $L_{Approx}(D)=C_{0}+A/D^{\alpha}$ for different context length. Previous work has succeeded in explaining this from an intrinsic space perspective, represented by (Sharma & Kaplan, 2022) as: $\alpha\approx c/d_{intrinsic}$ where $c=2\text{ or }4$ based on model property, and $d_{intrinsic}$ is the dimension of the intrinsic space of the dataset.
As assumed in Section 2.2.1, the Intrinsic Dimension should increase with $l$. Combined with previous results on $\alpha=c/dim(l)$, we have,
$L_{Approx}$ $=C_{0}+A(l)/D^{\alpha(l)},$ (5)
$\frac{\partial\alpha}{\partial l}$ $<0.$
This shows that longer context length would make it harder for the model to learn to approximate the Bayesian Model.
3 Deduction: Optimal Context Length and Training Dataset Size
In this section, we show a deduction of our theory presented in Section 2. We study the problem about a certain model trained on certain amount of training dataset $D$ with context length $l$, and validated on the validation set with the same context length $l$, where we want to know the impact of $l$ on validation loss.
As shown in Section 2.2, we can write Loss as:
$L_{CE}$ $=C_{0}+\frac{C}{l^{\gamma}}+\frac{A(l)}{D^{\alpha(l)}},$ (6)
In previous sections, we did not specifically discuss the relationship between $A$ and $l$. We consider $l$ where $\partial_{l}L_{CE}=0$ would give us an optimal $l$ with respect to $D$:
$\partial_{l}A=-A\ln D(-\partial_{l}\alpha)+\gamma C\frac{D^{\alpha}}{l^{\gamma+1}}=f(D,l).$ (7)
We find $\lim_{D\rightarrow 0}f=-\infty$ and $\lim_{D\rightarrow\infty}f=\infty$. This shows that for fixed $l$, no matter what $\partial_{l}A$ is, there exists some $D$ s.t. $\partial_{l}L_{CE}=0$.
We see that Bayes Risk decreases with $l$, Approximation Loss increases with $l$ but decreases with $D$; the balance between these two losses results in an optimal $l$ that increases with the optimal $D$.
We conduct experiments on a subset of OpenWebText with a sufficiently long context length. We trained GPT-2 on different context lengths with different amounts of training data, until the validation loss increases. We show our results in Figure 1 and Figure 5. Details for our experiment settings are presented in the Appendix F.
As shown both theoretically and experimentally, when training until overfitting on the training dataset and considering the perplexity on validation dataset, there does exist an optimal context length, beyond which even relevant long context would increase validation loss of pretraining Language Models. Such optimal context length would increase with training dataset size.
Figure 5: Openwebtext subset, Validation Loss vs. Context Length, for different dataset sizes. Different curves represent different amount of training data used. A more readable figure can be found in Figure 1, where the minimum validation loss reachable for each training dataset size is subtracted.
4 Proof of Concept with Synthetic Data
4.1 List of Points to prove
In this section, we conduct experiments on a synthetic dataset, explaining the Bayes Risk and related theories we proposed in Section 2.2. With this synthetic dataset, we would like to prove the following,
  • Point 1. Cross Entropy Loss is approximately linear with Intrinsic Dimension. Shown in Section 4.3.
  • Point 2. By measuring Eigen-value degradation in Intrinsic Space of well-trained models, one could obtain a valid measurement of the Intrinsic Dimension. Shown in Section 4.4.
  • Point 3. There exists an optimal context length for each training dataset size used, and such optimal context length increases with the amount of training dataset. Shown in Section 4.5.
4.2 Construction of Synthetic Data: the ‘position weighted multitask sparse parity’ dataset
In previous work, a common practice is to mask the leftmost tokens and leave $l$ tokens before the token-to-predict visible to Language Models, as shown in Figure 6. Although this may not show the impact of important tokens to final answer perplexity (e.g., it fails to show the importance of the second key info in Figure 6), this method aligns well with our setting of increasing context length.
Figure 6: An example of the ‘two needles in a haystack’ task, similar to those in (Levy et al., 2024). Left: the input to the Language Model, with key information and question visualized in blue. Right: perplexity of the answer token $\langle 8\rangle$ of LLaMa-3.1-8B (horizontal axis) vs. number of masked leftmost tokens (vertical axis). Although seeing both pieces of information are necessary to answer the question, perplexity rises dramatically only when the first piece of information is masked.
Although the next token to predict might depend on several pieces of key information, we see from Figure 6 that the first key token would raise model perplexity.
Inspired by this concept in Figure 6 and the ‘multitask sparse parity dataset previously studied in (Michaud et al., 2024; Barak et al., 2022), we propose the ‘position-weighted multitask sparse parity dataset. In detail, each input consists of $T$ ‘control bits and $L$ ‘context bits, each bit lies in $\{0,1\}$. Each subtask takes xor on two certain bits in the context bits, and the answer to some sample is the answer of the only activated subtask, as shown in Figure 7. We set tasks to have the same probability in the training and testing datasets. To emphasize the importance of context in different positions, for task $t$, we ensure $l(t)=max(context\_bit_{1}(t),context\_bit_{2}(t))$ follows $l(t)\approx 20/(1-t/50)^{1/1.2}$. This ensures,
$t(l)\approx 50-\frac{50}{(l/20)^{1.2}},$
which is the number of solvable tasks given a model of context length $l$.
Figure 7: An example of our synthetic data. The answer for Subtask 1,2,3 is $0\oplus 0=0,0\oplus 1=1$ and $1\oplus 1=0$ respectively, but since the thrid bit is $1$ for control bits, only Subtask 3 is activated and the final answer is 00. However, for a model of context length $7$, it cannot see the $9th$ bit required by subtask 3, making it unable to predict the answer correctly.
In usual cases, Intrinsic Dimension is equal or smaller than the number of subtasks given its definition to be the least dimension needed to capture useful information required to carry out all the subtasks in the system. We carefully design $min(context\_bit_{1}(t),context\_bit_{2}(t))$ (please refer to Appendix E for more details) to make the theoretical intrinsic dimension equaling to the number of solvable tasks.
Thus we have, for some model who presents $s$ bits of information (about the output logit) in one Intrinsic Dimension:
$ID(l,s)=t(l)/(s*T)\approx ID_{0}-C^{\prime}/l^{\gamma},$
in which, if we assume that the $1$ dimension in Intrinsic Space represents $1$ subtask (which has $1$ bits of information about the subtask and hence $1/T$ bits of information to the output logit), then $s=1/T$, thus $ID(l,1/T)=t(l)$.
4.3 Cross Entropy vs. Intrinsic Dimension vs. Context Length
We train a large enough MLP on data generated on the previous tasks, and evaluate our model on the validation dataset. During synthetic data generation, we make sure the training dataset and validation dataset do not overlap on any sample. We train until overfitting the training dataset. We assume $1$ dimension in Intrinsic Space can store information about $1$ subtask, hence we take $ID(l)=t(l)$ as its theoretical value here. After training, we obtain such results:
Let $f(x,C,C_{0},\gamma)=C_{0}-C/x^{\gamma}$ and $g(x,k,b)=k*x+b$.
The fitted results are:
  • ID vs. CL: $ID\approx f(CL,C,C_{0},\gamma)$,$C_{0}=-51.1\pm 1.0$, $C=-1.7*10^{3}\pm 0.3*10^{3}$, $\gamma=1.18\pm 0.06$, $R^{2}=0.9997$.
  • CE vs. CL: $ID\approx f(CL,C,C_{0},\gamma)$,$C_{0}=-0.015\pm 0.013$, $C=23.8\pm 4.3$, $\gamma=1.18\pm 0.06$, $R^{2}=0.9997$.
  • CE vs. ID: $CE\approx g(ID,k,b)$, $k=0.693\pm 1*10^{-5}$, $b\approx 0.0139\pm 4*10^{-7}$, $R^{2}=1-7*10^{-9}$.
As shown, we construct synthetic data such that $ID(l)=ID_{0}-C^{\prime}/l^{\gamma}$, and our measurements show $CE=C-C^{\prime}/l^{\gamma}$. More importantly, for the synthetic data example, Cross Entropy loss is almost perfectly linear with the Intrinsic Dimension as we defined previously. This validates Point 1: Cross Entropy Loss is approximately linear with Intrinsic Dimension; and we have also provided a construction to match the measured relationship $CE(l)\approx C_{0}+C/l^{\gamma}$.
4.4 Eigen Values in Intrinsic Space vs. Intrinsic Dimension
In this subsection, we train a model with a specialized architecture, allowing us to use the feature representation of a middle layer as a feature vector for input context bits, as shown in Figure 8.
Figure 8: Model trained on the proposed synthetic dataset; $\oplus$ represents feature concatenation. Only the first $l$ bits are used as input to context MLP when the context length is set to $l$. We conduct PCA on the Context Feature to analyze the intrinsic dimension of input context bits for various context lengths.
After training the model on data with different context length, we obtain the context feature representation and conduct PCA on it to study the Intrinsic Space of this model, presented in Figure 9.
From Figure 9 we find that: (1) the neural network would indeed learn the key information in the context bits. For models with different input context lengths, although their inner dimensions are the same ($80$), the representation of inputs in this inner space mainly lies in the first $ID$ dimensions, and the eigen values corresponding to other dimensions are very small; and (2) there exists such threshold $y^{*}$ that would estimate $ID$ for all context lengths accurately. In detail, using $y^{*}$ as the threshold to estimate $ID$ means that we find $idx\ s.t.\ rela\_eig(idx)>y^{*}$ and $rela\_eig(idx+1)<y^{*}$. We label a rectangle to show the possible $y^{*}$ for each context length and show that there exists $y^{*}$ to provide accurate estimations in Figure 9.
Figure 9: Illustration of relative eigen value vs. index, for models trained on different context length. Vertical lines: $x_{l}=ID(l)$ and $x_{l}^{\prime}=ID(l+1)$. For example, a context length $27$ has $16$ subtasks visible, corresponding to $16$ bits in Intrinsic Space. Assuming $1$ dimension in Intrinsic Space represents $1$ bit, the leftmost purple rectangle drawn means a range of $y_{threshold}$ that would provide an accurate estimation of $ID(27)=16$ for context length $27$. Obviously, there exists $y^{*}$ that would provide an estimation of $ID$ for all context lengths, as shown in the figure.
This provides evidence for Point 2 in Section 4.1: we can take some threshold $y^{*}$ to estimate the intrinsic dimension, by obtaining the maximum index of the relative eigen value such that the relative eigen value is larger than $y^{*}$, which would give accurate and consistent estimates.
The linear relationship between CE Loss and Intrinsic Dimension can also be explained from an Entropy in Intrinsic Space perspective, which is discussed in Appendix C.
4.5 Optimal Context Length and Training Dataset Size: Synthetic Dataset
To validate the idea that longer context lengths are suitable for larger datasets, we conduct experiments on our proposed synthetic dataset with a fixed size MLP. During our experiments, we change the dataset size and context length used, and obtain validation Binary Cross Entropy loss vs. Context Length, as shown in the lower figure in Figure 1. (We also plot $y=-\log(1/2)$ as the loss of random guess in Figure 1.)
From Figure 1, we see that models with longer context length need larger datasets to train, and training long-context models on small datasets could lead to worse performance or even failure to improve from the first step (performance equal to random guess). Moreover, for each curve that represents a certain size of the training dataset used, we observe an optimal horizon beyond which validation loss would increase. This critical point moves right (optimal context length increases) with the increase of the training dataset size. This proves point 3 in Section 4.1.
5 Related Work
5.1 Enlarging Context Length for LMs
Previous work has made attempts to enlarge the context length of Language Models. Work represented by RoPE(Su et al., 2023) uses rotary positional embedding to support generalizing LMs to longer context in inference compared to the training process. These work uses modified positional embeddings to model the relative position dependency in attention mechanism.
There is also work about enhancing long context understanding and exploring Scaling Laws for context length(Xiong et al., 2024). These work utilize an adjusted pretraining and instruction-finetuning process with more long-context data to enhance the models’ ability on long contexts.
Other work modifying architectures has also been proposed to enhance long context modeling or to simplify deployment of long context LLMs. For example, (Tang et al., 2024) proposes a training-free RazorAttention algorithm to largely compress the KV cache while maintaining performance unchanged.
Architectures and inference methods have been proposed to reduce inference time and memory cost for Language Models, represented by a series of linear transformer or RNN-based methods.(Katharopoulos et al., 2020; Gu & Dao, 2024; Sun et al., 2024). These methods, largely reducing the computational cost and memory usage for long input contexts, have displayed margin ahead of traditional attention-based langauge models for long context inference.
Currently a common practice to train very large Language Models supporting long context is to use pretrain the model with shorter contexts, then finetune them with longer contexts, as presented in tech reports of LLaMa-3(Grattafiori et al., 2024) and DeepSeek-v3(DeepSeek-AI et al., 2024).
5.2 Irrelevant Long Context hurts performance of LMs
Besides context length scaling with relevant contexts, previous researches have studied how LLMs perform for long irrelevant contexts. As an example, (Levy et al., 2024) studies the performance of current LLMs on an adjusted version of ‘needle in a haystack task, where two pieces of key information are embedded into a long text corpora and a question related to both is asked, similar to that presented in Figure 6. The conclusion of these work is that LLMs would perform worse when there is too much irrelevant information.
5.3 Long Context in another field: Time Series Forecasting
Context length, representing the length of input context, is not unique to Nature Language. For time series forecasting, where machine learning plays an important row, there is also work discussing the impact of context length, represented by (Shi et al., 2024). These investigations find that there exists an optimal look-back horizon, which increases with dataset size. However, time series datasets are relatively small compared to NLP datasets, and thus whether this conclusion holds on NLP remains an open problem for this work to study.
5.4 Related Theories for Scaling Laws
Since the discovery of Scaling Laws for Large Language Models (Kaplan et al., 2020) or even earlier, there has been theoretical work trying to explain why model performance could benefit from more data points and more model parameters. For exmaple, (Sharma & Kaplan, 2022) studies the dataset and model scaling from the data manifold perspective.
Specially for Language Models, there is also previous work proposing all kinds of theoretical models. For example, (Michaud et al., 2024) proposes a feature-quant based theory; (Aghajanyan et al., 2020) views the effect of fine-tuning from the intrinsic dimension perspective; (Havrilla & Liao, 2024) proposes to understand scaling with intrinsic dimensions.
6 Conclusion and Discussions
6.1 Conclusion
In this work, we discuss the impact of context length on language modeling, especially Bayes risk and approximate loss, from both a theoretical and experimental perspective.
In Section 2, we propose assumptions related to the relationship between CE Loss, intrinsic dimension and context length. We derive a linear relation between CE loss and Intrinsic Dimension, and study the impact of context length to intrinsic dimension. We further investigate the relationship between intrinsic dimension, context length, and Entropy in intrinsic space in the appendix C from an Entropy perspective.
We also conduct experiments on real data (Section 2, Section 3) and synthetic data (Section 4), on measuring Intrinsic Dimension and on the relationship between Cross Entropy Loss (Bayes Risk and Approximation Loss), Context Length and Intrinsic Dimension.
As a correlation of our theory, for the training-till-overfitting setting, there exists an optimal context length that increases with dataset size in the pretraining process. This is also validated in Section 3 and Section 4.5.
We hope our work may provide insight for future work about long context Language Models, or about Physics for Language Models.
6.2 Limitations and Future Work
In Section 3 we mainly discuss the case for pretraining with long context, but common practice is to train with shorter contexts and then do long-context-training with longer contexts (DeepSeek-AI et al., 2024). We leave these to future work.
Our theory starting from Intrinsic Dimension only holds with assumptions in Section 2; and in Appendix C we use the perspective of Entropy and possible states number in Intrinsic Space to (partially) explain our assumptions and measurements w.r.t. Intrinsic Dimension. We hope future work may try to propose even more fundamental theories to explain our Intrinsic Dimension measurements and assumptions.
Though not discussed in detail in this paper, there are problems that could be potentially explained by our theoretical framework. For example, in the measurement conducted in this work, the Intrinsic Dimension with increasing context length converges to a finite point, which could potentially explain why RNN models(Gu & Dao, 2024; Peng et al., 2023; Sun et al., 2024) with limited hidden state can be good Language Models.
Statement for Possible Impact
This paper presents work whose goal is to advance the field of Machine Learning (or more specifically, about understanding the impact of long contexts in Language Modeling). There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Definition of Cross-Entropy loss used in this work
It is well-known that
$H_{org}(P,Q)=\sum_{x}$ $-P(x)\log Q(x)$ (8)
$=\sum_{x}$ $-P(x_{-\infty:0})P(x_{0}|x_{-\infty:0})$
$*\log\{Q(x_{0}|x_{-\infty:0})Q(x_{-\infty:0})\},$
where $x_{a:b}$ denotes $x_{a},x_{a+1},\ldots,x_{b-1}$. Although this should be the definition of $H(P,Q)$, it is common practice to calculate perplexity of Language Models with its input as GT lables (e.g. in technical report of LLaMa-3(Grattafiori et al., 2024)), in other words,
$H_{exp}(P,Q)$ (9)
$=$ $\sum_{x}-P(x)\log Q(x)$
$=$ $\sum_{x}-P(x_{-\infty:1})\log\mathbf{\{}Q(x_{0}|x_{-\infty:0})\mathbf{P(x_{-\infty:0})}\mathbf{\}}$
$=$ $\text{Const}+E_{x_{-\infty:0}}\sum_{x_{0}}-P(x_{0}|x_{-\infty:0})\log Q(x_{0}|x_{-\infty:0}).$
Therefore, in this work we use:
$H(P,Q)$ (10)
$=$ $\text{E}_{x_{-\infty:0}}[\sum_{x_{0}}-P(x_{0}|x_{-\infty:0})\log Q(x_{0}|x_{-\infty:0})]$
as the definition of Cross-Entropy loss, and $P(x_{0}|x_{-\infty:0})$, $Q(x_{0}|x_{-\infty:0})$ as the definition of Nature Language distribution and Language Model distribution, respectively.
Appendix B Detailed Derivation for KL distance in Section 2.2
Based on the definition in Appendix A, we derive $D_{KL}=(dim(\infty)-dim(l))*S$ here.
$D_{KL}(P,P_{l})$ (11)
$=E_{x_{-\infty:0}}\sum_{x_{0}}D_{KL,x_{0}}(P(x_{0}|x_{-\infty,0}),P_{l}(x_{0}|x_{-\infty,0}))$
$=(dim(\infty)-dim(l))*S$
This is because we assume that each dimension in the intrinsic space can store $S$ bits of information.
Another intuitive example is: assuming that the vocab is an integer from 00 to $12^{dim(\infty)*S}-$ 1. assuming $P(x_{0}|x_{-\infty:0})=\delta_{x_{0},y}$, that is, the next token given $x_{-\infty:0}$ is sure to be $y$. $y$. For $P_{l}(x_{0}|x_{-\infty:0})$, the first $dim(l)S$ digits of the integer (in binary representation) are known, but the remaining $(dim(\infty)-dim(l))S$ digits are unknown, making a guess in these numbers yield $P_{l}(x_{0}|x_{-\infty:0})=1/2^{S*(dim(\infty)-dim(l))}$. Thus, $D_{KL,x_{0}}(P(x_{0}|x_{-\infty,0}),P_{l}(x_{0}|x_{-\infty,0}))=1*\log 1/(1/2^{S*(dim(\infty)-dim(l))})=S*(dim(\infty)-dim(l))$.
Appendix C Entropy perspective: hyper-volume in Intrinsic Space
C.1 Bayes Risk from an Entropy-in-Intrinsic-Space perspective
Here we derive similar results as in Section 2.2, but not from an Intrinsic-Dimension perspective, rather we derive it from an Entropy perspective.
  • Assumption 1. Entropy of the Bayes Model $\lim_{l\rightarrow\infty}H_{ntp}(l)=H_{ntp}(\infty)$ is finite, which is the Entropy of next token prediction of language itself.
  • Assumption 2. $\forall l_{1},l_{2}\text{ such that }l_{1}<l_{2}$, $H_{ntp}(l_{1})<H_{ntp}(l_{2})$. This is because a longer context contains more information about the next possible token.
To simplify deduction, we further assume that,
  • Assumption 3. ${D_{KL}}_{,ntp}(P,P_{l})\approx H_{ntp}(P)-H_{ntp}(P_{l})+Const$. This means that the KL divergence between the Bayes LM of infinite context length and the Bayes LM of context length $l$ can be written approximately in the form of entropy difference.
  • Assumption 4. The Entropy w.r.t. Next Token Prediction, i.e., $H_{ntp}(P_{l})$, is linear with the Entropy in the Intrinsic Space of the Bayes Model, i.e., $H_{IS}$.
To establish a relationship between Cross Entropy and Intrinsic Space, we run LLaMa-3-8b on a subset of the Openwebtext dataset and obtain the feature of the last token as the feature representation, or Intrinsic Space of the approximated Bayes Model, as shown in the left part of Figure 10. We see that the model with larger context length tends to have larger relative eigenvalues in intrinsic space, thus containing more information.
Figure 10: Left: Relative Eigen Value Measured for the last token, for LLaMa-3.1-8B on a subset of OpenWebText. Right: relative increment of relative eigenvalues (for different context lengths measured). We can see that the relative eigenvalues approximately increase at a same scale.
Entropy of a system can be defined as $H=\log\Omega$ where $\Omega$ is the possible number of states of the system. Similar to the calculation process of Entropy in Statistical Physics(Landau & Lifshitz, 1980), we define Entropy of Intrinsic Space as:
$H_{IS}$
$=$ $\log\Omega$
$=$ $\log V/h^{dim(V)}\text{ where $V$ is the volume in intrinsic space}$
$=$ $\sum_{i}\log rel\_eigval_{i}/h$
$=$ $\sum_{i}\log rel\_eigval_{i}+Const$
Here $h$ is the ‘plank constant’, meaning that one state corresponds to a unit hyper-volume of $h^{dim}$ in the Intrinsic Space. A different value of $h$ would only add a constant to $H_{IS}$ and would not affect change in Entropy. Thus, we use $\sum_{i}\log rel\_eigval_{i}$ as Entropy in Intrinsic Space. Moreover, experiments in Figure 10 show that relative eigenvalues increase in the same ratio as the context length increases. This means the number of states in any subspace would scale proportionally with respect to the number of states in the whole space, that is: $H_{subspace}=H_{IS}*dim(subspace)/dim(IS)$, hence Entropy of subspaces are linear with respect to each other.
Though related, Entropy in Intrinsic Space does not equal to Entropy in the next token prediction task. From the probability perspective, let $dec(x)$ be the next decoded token for some point $x$ in the intrinsic space, we have: $H_{IS}=\sum_{x\in IS}-P(x)\log P(x)$, while $H_{ntp}=-\sum_{v\in vocab}P(v)\log P(v)$ where $P(v)=\sum_{x\in IS,dec(x)=v}P(x)$. $H_{ntp}$ is a coarse-grained Entropy compared to $H_{IS}$. $H_{IS}$ contains important information on previous tokens that are important for the prediction of future tokens, while $H_{ntp}$ is related only to the next token.
Experiments in Figure 11 show that, no matter what subspace we use, the Cross Entropy Loss usually follows a linear relationship with the Entropy we measured in the subspace, supporting the claim that the next token prediction task likely lies in some subspace of the Intrinsic Space, or (statistically) its Entropy should be some weighted average of Entropy of several subspaces of similar dimension.. This also suggests that $H_{ntp}$ is approximately linear with $H_{IS}$, which validates our previous assumptions and claims.
Figure 11: Upper-left, Upper-right, Bottom-left: Cross Entropy Loss vs. $\sum_{i\leq N}\log$ rel_eig_val; Bottom-right: correlation between minus CE loss and $\sum_{i\leq N}\log$ rel_eig_val. All experiments are for LLaMa-3.1-8B on a subset of OpenWebText. From the first three figure, we see CE loss is linear with the Entropy of certain subspaces. From the bottom-right figure, we see that Entropy measured in different subspaces are highly correlated ($corr>0.97$), which are also highly correlated with the CE loss for Next Token Prediction.
C.2 Bridging the gap between Intrinsic Dimension explanation and Entropy explanation
Here, starting from previous assumptions and measurements w.r.t. Entropy in Intrinsic Space, we explain why CE is linear w.r.t. Intrinsic Dimension measured in Section 2.2, for $idx>500$. We see in Figure 3 that for $idx>500$, the eigenvalues mainly follow an exponential decay:
$releigval_{ctl,\ idx}=releigval_{ctl,\ 0}*\exp\{-\alpha_{ctl}*idx\}\text{, forcertain context length}$
We also now from the previous result that for different context lengths, the relative eigenvalues increase almost in the same proportion, especially for $idx>1000$. That is, $\alpha_{ctl}=\alpha$, and $releigval_{ctl,\ idx}=\gamma(ctl)*releigval_{\infty,idx}$.
For the subspace for the next token prediction task, the entropy should be proportional to log of volume in the subspace; hence it should be proportional to $m*\log\gamma(ctl)$ where $m$ is the dimension of this exact subspace. That is:
$H_{subspace}(ctl)=m\log\gamma(ctl)+Const$ (12)
For some certain threshold $thres$, we use, for some context length, the measured intrinsic dimension is:
$releigval_{\infty,0}*\gamma*\exp\{-\alpha*dim(ctl,thres)\}=thres,$
hence $\gamma(ctl)=thres/releigval_{\infty,0}*\exp\{\alpha*dim(ctl,thres)\}$. Plugging this into Equation (12) we have:
$CE=-H_{subspace}(ctl)+Const=-m\alpha*dim(ctl,thres)+Const(thres).$ (13)
Thus, we derive our assumptions in Section 2, where $s=m\alpha$. Equation (13) can also be validated in the fourth part of Figure 4, where the measured Intrinsic Dimensions (for $idx\geq 500$) are measured in the exponential decay area, and share similar slopes w.r.t. CE Loss.
C.3 Synthetic dataset: Entropy in Intrinsic Space, and Entropy for output layer
For our synthetic dataset, if we view the Context Feature Vector shown in Figure 8 as the feature in the Intrinsic Space, then the best strategy for the context encoder is to generate the answer for all subtasks (it can see) in the Intrinsic Space (since it cannot see the task bits). This would lead to an Entropy of $H_{IS}=T\log 2$ in the Intrinsic Space.
The entropy of the output layer is, however, $H_{output}=\log 2$ since the answer bits $0,1$ have the same probability. In this way, the answer of the output layer actually corresponds to one dimension in the Intrinsic Space, which should be the exact dimension at which the answer of the current task is stored. Therefore, $H_{output}=1/T*H_{IS}$, which explains why the Entropy for output logits is linear to the Entropy for Intrinsic Space.
C.4 New Synthetic dataset: further experiments for Entropy
In the main paper, we use $T<ctl$, where the task number is smaller than the context bits. Here, we further study the case where $T>ctl$. Moreover, instead of letting tasks show at the same frequency, we use weighted frequency proportional to $max(bit_{1},bit_{2})^{-1.2}$ as the weights of tasks.
Here we train a single MLP (instead of multiple MLPs) as Bayes Model for different context lengths. The architecture is like:
Figure 12: New MLP trained with mask bits. For the example shown, the context length is $7$ and we mask the other task bits to be $0.5$.
We use $100$ tasks and $60$ task bits. From $11th$ to the $60th$ bit, each bit corresponds to the max bit of two tasks: that is, #$\#Task|_{max(bit_{1},bit_{2})=i}=2,\forall i\in\{11,12,\ldots,60\}$.
During training, $50\%$ of the samples are unmasked, while for the other $50\%$ samples, we mask the last $X$ task bits to be $0.5$, where $X$ is a random int from $60-10$ to $60-60$. This ensures our model to be able to handle mask bits, and also ensures it can learn uncommon tasks (relying on context bits that are at the end of the context bits) well.
The context bits MLP has hidden size $400$, and the prediction MLP has hidden size $300$. The linear embedding and the context feature have the same dimension $200$, and are added before being sent to the predict MLP. LeakyRELU with leaky hyperparameter $31e-$ 3 is used as an activation function for both MLPs. We train the model on a training set of $10000000$ and a validation set of size $1000000$, for $125$ epochs (and an early stopping setting of $25$ epochs, though the training process did not trigger early stopping).
To make sure that the trained model can be used to approximate the Bayes Model, we compare the model’s loss on validation set with context $ctl$ with the calculated minimum possible CE Loss for the task:
$MinCELoss(ctl)=\frac{\sum_{\text{task s.t. $max(bit_{1},bit_{2})>ctl$}}P(task)*\log 2}{\sum_{task}P(task)}$
And we obtain such a result:
Context Length Model CE Loss Minimum CE Loss Calculated
15 0.531 0.516
17 0.478 0.464
20 0.408 0.399
23 0.353 0.344
25 0.320 0.312
28 0.276 0.269
30 0.250 0.243
35 0.193 0.186
40 0.154 0.139
50 0.082 0.061
60 0.051 0.0
Table 1: Comparison between trained model and Bayes Model for Synthetic Data
We can see from Table 1 that the model is not too different from the Bayes Model: the BCE Loss only differs by around $0.02$. Thus, we can use the middle-layer-representation (shown as context feature in Figure 12) as the feature in Intrinsic Space to approximate the Bayes Model for $17\leq ctl\leq 50$.
Figure 13: Eigen value and CE results measured on trained model for Synthetic Dataset in this section. Left: relative eigen value vs. Index. Middle: Model CE Loss vs. $\sum_{i}\log rel\_eigval_{i}$ and Theoretical Min CE Loss vs. $\sum_{i}\log rel\_eigval_{i}$. Right: Correlation between minus Experimentally measured model CE Loss, minus Theoretically Minimum CE Loss, and $\sum_{i\leq N}\log rel\_eigval_{i}$ for different $N$s. All correlation factors are larger than $0.97$.
We show our results in Figure 13. We see from Figure 13 that: (1) Larger context length contains more information, hence eigen values in Intrinsic Space degrades slowlier (left figure); (2) the model approximates the theoretical Bayes Model well (as the green points in the middle figure is very close to the orange ones) (middle figure); (3) CE Loss follows a very good linear relationship with sum of log eigenvalues of the first $N$ dimensions for $N\geq 70$ in the Intrinsic Space (right figure), where the case $N=200$ (all eigen values) are also shown in the middle figure.
Appendix D More experiments of LLaMa on another dataset
According to the technical report of LLaMa 3(Grattafiori et al., 2024), the text corpora with number of ‘dirty words’ beyond certain threshold would be filtered out, as proposed in (Raffel et al., 2023). We collect some text corpora online which include forbidden words defined in (Raffel et al., 2023), as text corpora unseen by LLaMa 3. By conducting experiments on it we obtain results similar to Openwebtext subset.
Figure 14: Cross Entropy Loss vs. Context Length, with log scale. We see that $y=C_{0}+C/x^{\gamma}$ fits this curve well.
According to Figure 14, we see that $CE=C_{0}+C/l^{\gamma}$ approximates well for text corpora that are sure not to be seen by the model.
Appendix E Construction of the Synthetic dataset
If the subtasks defined in Section 4.2 is independent with each other, and number of subtasks visible is smaller than number of context bits visible, then the intrinsic dimension should equal to the number of subtasks: we need $T$ bits to store the required information of these subtasks. However, in special cases when different tasks are too dependent with each other, then the number of bits needed to represent the answer is $4$ instead of $T=5$; for example: if we have $4$ context bits, and $T=5$ tasks dependent of context bits $\{(1st,2nd),(1st,3rd),(2nd,3rd),(3rd,4th),(1st,4th)\}$ respectively, then the Intrinsic Dimension should be $4$ instead of $5$. This would make Intrinsic Dimension less than the number of tasks defined.
To avoid this we carefully tune the bits dependency of our tasks defined, and list them as follows:
[…]
$(23,6),(23,7),(24,8),(24,9),(25,10),(25,11),$
$(26,12),(26,13),(27,14),(27,15),(28,16),$
$(29,17),(30,18),(30,19),(31,20),(32,1),(33,3),$
$(34,6),(35,8),(36,10),(37,12),(39,14),(40,16),$
$(42,38),(43,41),(45,44),(47,46),(50,48),(52,51),$
$(55,54),(58,57),(62,61),(66,65),(71,70),(77,76),$
$(84,83),(93,92),(103,102),(118,87),(137,25),$
(165,20),(209,34),(293,128),(522,353)(165,20),(209,34),(293,128),(522,353)\}( 165 , 20 ) , ( 209 , 34 ) , ( 293 , 128 ) , ( 522 , 353 ) }
One can check that the task defined here has an intrinsic dimension for any model of context length $l\geq 23$ equal to the number of subtasks visible for that specific context length.
Appendix F Experiment Settings
F.1 Natural Language Data
F.1.1 Optimal Context Length Experiments
We use aanogpt(Karpathy, 2022) and train a model with GPT-2(Radford et al., 2019) architecture on a subset of OpenwebText dataset. Our model is the same with GPT-2-124M ($12$-head transformers, $768$-dim feature vector) except that it uses half the transformers layer size ($12\rightarrow 6$) to reduce GPU memory for long contexts. For training, we use the AdamW(Loshchilov & Hutter, 2019) optimizer, an equivalent batch size of $480$ learning rate of $46e-$ 4, weight decay of $11e-$ 1, $2000$ warm-up iterations and $600000$ lr-decaying iterations. We train until overfitting, that is, we validate the validation loss of our model on the validation set every $45$ steps, and choose the best iteration as the min validation loss measured for a single run.
We train the model on a subset of OpenWebText. To be specific, we first select text corpora with context length beyond specific limits larger than the maximum training context length from OpenWebText, then split into Training set and Validation Set. The training set we used to train the models have $12M,25M,50M$ tokens respectively, and the validation set has $134M$ tokens.
Each point in the corresponding figure in Figure 1 and Figure 5 is produced with $8$ AMD MI-250X GPUs (which are similar in performance to Nvidia A100 gpus) trained for $2$ days.
F.1.2 Intrinsic Dimension Experiments
We select long enough text corpora from the Openwebtext dataset. Then, following previous practice(Cheng et al., 2023), we conduct experiments with LLaMa-3.1-8b on $10000$ samples of this subset. We extract the feature representation of the last token in the last layer, as the Intrinsic Representation of samples.
Conducting all intrinsic dimension measurements cost up to around $100$ gpu hours for MI-250X gpus.
F.2 Synthetic Dataset
F.2.1 Settings in Section 4.3 and 4.5
We use a validation dataset size of $2000000$. Training dataset size is $2000000$ in Section 4.3, and varies as shown in Figure 1 in Section 4.5.
We use a large enough MLP on large enough datasets. To be specific, our MLP has four linear layers and two leaky-relu activation layers. The input has shape $context\_length+taskbits\_length$, and output has shape $1$. The hidden layer is of dimension $400$ and middle layer is of dimension $200$. During training, we use the Adam(Kingma & Ba, 2017) Optimizer, batch size is $10000$, learning rate is set to $31e-$ 3, weight decay is $41e-$ 4 and we train until validation loss increases for $25$ epochs, with a maximum epoch number of $200$.
For Section 4.3, these settings enable the model to be fully trained on enough data, to be viewed as a Bayes Model.
F.2.2 Settings in Section 4.4
For models shown in Figure 8, we use two-layer mlp for both context feature encoding and logit decoding. The hidden sizes are $400$ and $200$, respectively, and the context feature dimension is set to $80$. The training rests are similar to the setting in Section 4.3.
Similar training settings are used for training Synthetic Data for Appendix C.
The experiments for Synthetic Data take around $10$ hours on an Nvidia RTX-3090 gpu.